# Mass Shootings, Community Mobility, and the Relocation of Economic Activity

### v1

@@ 4.1 Mobility Patterns after Mass Shooting
Figure 2(a) plots event study estimates obtained from model (1) for two distance bands split across median distance from the shooting site. POIs near shooting sites experience statistically significant declines in weekly visits for most of the period. We do see some degree of recovery after week 5, but it does not persist in the later weeks. The pattern is robust to an alternative specification using four distance bands instead of two (online Appendix B), though the effect is less precisely estimated for the 1st quartile distance band, which may be due to the presence of authorities, visitors, or the media near the shooting sites.
@@
POIs above the median distance, on the other hand, show an increase in visits starting a few weeks after the event. Although the POIs also exhibit a decrease in weekly visits until week 5 (statistically insignificant), visits return to the pre-event level in week 6 and remain above the baseline for the remainder of the period. Two potential factors may explain the increase in weekly visits to these POIs. First, the resumption of previously postponed economic activities by individuals who normally use those POIs. Second, the relocation of activities to further away places by individuals who, prior to the event, typically used the POIs near the shooting site. The former implies only a temporal displacement of activities, but the latter suggests both temporal and spatial displacements. Although the resumption of postponed activities can account for the recovery observed between weeks 5 and 7, it does not fully explain why the recovery goes beyond the pre-event level and the increased visits persist for more than three months. Another important observation in Figure 2(a) is that, after week 5, the zero horizontal line often passes through the middle of the gap between the two plots, indicating the net change in weekly visits across all regions is often close to zero.
@@
Together, these patterns suggest some degree of substitution or relocation of economic activities across the regions. That is, visits lost from the POIs in close proximity to the crime scene are redistributed to the POIs far from it. [^12] This relocation mechanism is further supported by the analysis in Section 4.2, where we examine the impact by POI type and POI substitutability. This substitution effect, however, does not kick in immediately after the shooting. Instead, a few weeks go by before a significant gap is observed in the post-shooting trends of weekly visits between the regions. This may be due to the difficulty of substitution in the presence of search and switching costs (Klemperer, 1987; Wilson, 2012; Weizsäcker, 1984). Therefore, the relocation of economic activities implies friction costs or efficiency losses due to a deviation from the optimal choices, which, in turn, could affect more macro-level economic outcomes such as wage, unemployment rate, and housing price, among others (a la Brodeur & Yousaf, 2022). [^13] Compared to mere temporal displacement, the relocation of activities can result in more enduring and persistent social costs since the new location is likely sub-optimal compared to the previous one.
@@
Note that the changes in weekly visits in Figure 2(a) are “per POI”; therefore, the aggregate impact can be substantial. Since there are over 150 thousand POIs in the data, the overall impact will be the product of the estimated change per POI and the number of POIs in a given distance band. For instance, the reduction of about 6 visits per POI in week 5 for those POIs below the median distance will amount to an overall decrease of almost half a million weekly visits. The median value of nonzero weekly visits in the dataset during the week prior to the shootings (the baseline week) is 20, and thus a decrease of 6 visits constitutes a 30% decline relative to the median.
@@
The differential mobility patterns in the two regions persist over time, resulting in a significant gap in weekly visits. Figure 2(b) plots the DiD estimates for POIs below the median distance (i.e., nearby POIs) with the reference group being the POIs above the median distance (i.e., far away POIs). The graph thus directly quantifies the gap in weekly visits over time between the two regions and its significance. The gap becomes statistically significant after week 4 and increases to about 8 visits in week 17. The same pattern is observed from an alternative specification with four distance bands, although estimates for first-distance-quartile POIs lack precision (see online Appendix B).
@@
Our results in this section are robust to the thresholds used to define distance bands. Online Appendix B.2 provides some nonparametric results based on splines (piece-wise polynomials), where we estimate the evolution of weekly visits as a function of “continuous” distance. The patterns are consistent with those reported in this section. The spline results also support setting 5 miles as the cutoff for sample selection, as estimates approach zero.
@@ 4.2.1 Impact by POI Type
Figure 3: Change In Weekly Visits by POI type
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. ”Near”/”Far” refers to POIs below/above median distance from the shooting location.
Figure 3 reports the changes in visits and corresponding 95% confidence intervals for various categories in the week immediately after the event (Week 0) and approximately two months after the shooting (Week 8). [^14] Online Appendix D.1 includes the results for Week 2, 4, 6. The estimates indicate that decreases in visits are immediate and quite general across a variety of POI types near the shooting sites, pointing to widespread economic costs of mass shootings. Notably, for most categories, the increase in visits to farther-away POIs closely mirrors the decrease to nearby POIs, supporting the relocation hypothesis. People start favoring POIs farther away from the shooting site.
Heterogeneity of the impact across POI types is also notable. Recreation, education, temporary accommodation, food/drink stores exhibit larger changes, although their confidence intervals are also wider. The large and significant effect on education immediately after the shootings deserves a look. Disruption of educational activities not only hampers student learning, but also negatively impacts parents’ economic activities such as reduced work hours and productivity (Garcia & Cowan, 2024). Therefore, this large effect on education suggests a substantial and broad social cost.
@@ 4.2.2 Impact by POI Substitutability
The relative size of the changes in weekly visits across different POI types in Figure 3 suggests that ease of substitution might play a key role in the observed patterns. It is relatively easy to change the recreational spaces you frequent and thus we see a significant shift in visits from nearby places to far-away places for recreation. In contrast, it is difficult to find a new day care, which accounts for the zero effect on daycares two months after the shooting.
To generalize the idea, we evaluate how the substitutability of POIs is associated with the magnitude of the impact of mass shooting. We perform two DiD regressions, similar to that in Figure 2(b) (i.e., calculating the changes in weekly visits to nearby POIs using faraway POIs as the reference group), for two different groups : high-substitutability and low-substitutability POIs. Substitutability is defined in two distinct ways, as shown in Table 4(a). The Common Categories measure of substitutability accounts for the nature of business based on our own experience but can be subjective. The Competition Level measure takes into account the level of competition in a community but may be endogenous since the number of competitors or businesses is determined by local demand or differential preference among consumers. [^15] That is, even though there are many competing locations, it may be a result of many different tastes among consumers, so substitution can still be limited. Each measure has its pros and cons, so the results in this section should be used to gain insights, rather than quantify the difference.
Figures 4(b) and 4(c) report DiD estimates based on the Common Categories and the Competition Level measures, respectively. In both cases, there are distinct patterns in the evolution of foot traffic over time, with the redistribution of visits from nearby to farther-away POIs being more pronounced for highly substitutable POIs toward the end of the period. Due to the smaller sample size for each DiD regression, these differences are not statistically significant. Nevertheless, the differential patterns are evident in these plots and the patterns are robust to an alternative specification using 4-quartile distance bands in online Appendix D, supporting substitution as the key mechanism driving the mobility changes in the data.
@@ 4.3 Degree of Traumatic Stress and Magnitude of Impact
The previous findings indicate that, after a mass shooting, people relocate their activities to far-away POIs. Given that the likelihood of another mass shooting occurring at the exact same location is nearly zero, the behavioral change should be understood in light of the psychological effects mass shootings have on communities. It has been well documented in the psychology and health literatures that traumatic events take a toll not only on survivors but also on their communities. Even those not directly affected by the event can develop post-traumatic stress symptoms, which is especially likely in cases of mass violence (Galea et al., 2003) and for those individuals who are physically closer to the trauma (Furr et al., 2010; Matt & Vasquez, 2008; May & Wisco, 2016).
A variety of studies find that fear of crime leads to considerable deterioration of both mental health and physical functioning (Green et al., 2002; Guite et al., 2006; Stafford et al., 2007). Wozniak et al. (2020) report that, after a 2014 university shooting, students, faculty, and staff expressed to both willingly and unwillingly reminisce on the traumatic event, and that such reminiscing was connected to the development of post-traumatic stress symptoms. Similarly, Sharkey & Shen (2021) analyze daily survey data and find that mass shootings cause people nearby to experience substantial shifts toward negative emotions. Furthermore, findings in the Economics literature confirm the existence of mental effects of crime, and add that, when people perceive an area to be unsafe, their willingness to pay decreases, and they attempt to adapt their behavior to avoid those areas (Dustmann & Fasani, 2014; Buonanno et al., 2013; Braakman, 2012).
@@
To examine how the degree of traumatic stress in our context influences our results, we split our sample into high- and low- traumatic stress groups based on three factors: preexisting crime rate, media coverage, and pre-event visit density. We then analyze how these factors affect the magnitude of the impact of mass shooting. Table 5(a) summarizes how we split the data and predictions based on the literature. The analyses in this section are similar to those in Figures 4(b) and 4(c), and supplemental plots are provided in online Appendix E.
@@
4.3.1 Preexisting Crime Rates
Two theoretical frameworks from psychology inform our predictions regarding how mass shootings may have different effects depending on the preexisting levels of crime in the affected area. The first is stress inoculation, which states that exposure to smaller instances of stress can promote resilience to future stressful episodes (Ashokan et al., 2016; Gerber et al., 2018). For example, Amir & Sol (1999) find that students in an Israeli university experienced decreased levels of distress when facing a traumatic event if they had previously experienced other traumatic events. The second, desensitization, consists of emotional numbing resulting from repeated exposure to chronic community violence (Kennedy & Ceballo, 2016; Gaylord-Harden et al., 2011; Mrug et al., 2008; Ng-Mak et al., 2004). It follows that the impact of mass shootings is expected to be greater when it occurs in a previously safer area, where the event would strike people who are less habituated to violent crime.
We match each mass shooting with the homicide rate of the county where the shooting occurred in the year prior to the event. Homicide statistics are sourced from County Health Rankings and Roadmaps [^16] https://www.countyhealthrankings.org. Then, we classify the shooting areas into two groups: high-crime areas, defined as those with above-median crime rates, and low-crime areas, including all other locations.
Figure 5(b) plots the evolution of weekly visits for high- and low-crime areas. Over time, the impact of mass shootings on mobility becomes stronger when they occur in previously safer areas. That is, when communities are less habituated to violence, their reaction is stronger, shifting activities more substantially from nearby POIs to those farther away.
4.3.2 Extent of Media Coverage
As discussed previously, fear of crime leads to the deterioration of mental health outcomes. Moreover, evidence suggests that its psychological impact is driven primarily by perceived risk, not by actual recorded crimes (Pearson & Breetzke, 2014; Braakman, 2012; Hamermesh, 1999; Hale, 1996). The media, then, would play an important role through its influence on public perception. According to News Media and Crime Perceptions: Evidence from a Natural Experiment (2018), who use a staggered introduction of digital TV in Italy as a natural experiment, reduced exposure to crime news channels causes people to decrease their concerns about crime. Similarly, Velasquez et al. (2020) find that spikes in crime reporting increase people’s perceptions about the prevalence of crime, even when actual crime rates decline. It follows that, all else equal, more extensive media coverage on mass shootings is likely associated with greater psychological distress at the community level, and thus stronger impacts on community mobility patterns.
@@ 4.3.2 Extent of Media Coverage
Shooting events are then divided in two groups: those receiving above-median media coverage (high coverage) and those below this threshold (low coverage). Figure 5(c) plots the evolution of weekly visits for high- and low- media coverage shootings. Consistent with our prediction, more extensively publicized events tend to result in greater shifts of visits from nearby POIs to farther-away POIs.
@@ 4.3.3 Visitor Density
Fear of crime is associated with an individual’s perception of vulnerability – that is, the self-assessed likelihood of victimization. Though such perceptions might not accurately reflect past personal experience with victimization, they can translate into economically-relevant behavioral changes, such as the avoidance of certain physical areas or situations where the perceived risk of victimization is high (Hale, 1996; Garofalo, 1981; Whitley & Prince, 2005). Furthermore, the decision to avoid a physical environment depends on the individual’s assessment of potential risks, even in absence of direct connection between the location and specific instances of crime. (Marquet et al., 2020; Navarrete-Hernandez et al., 2023).
@@
The results are consistent with the prediction that communities respond more intensely for the locations where larger gatherings of people are expected, likely in an effort, conscious or not, to avoid situations that are perceived as conducive to victimization. POIs with low baseline visit density exhibit little to no change in visits after a mass shooting, with very narrow confidence intervals. The median number of nonzero weekly visits prior to the shootings among all POIs in the sample is 20, suggesting that the POIs in this category are generally very small and, as a result, do not show meaningful variations in visits. [^18] This may indicate that the effects of mass shootings presented elsewhere in the paper can be seen as a lower bound, as they include POIs that are expected to show little change in visits due to the nature of the business.
5 Conclusion and Discussion
In the current reality, where the tragic recurrence of mass shootings brings the issue into the spotlight of political debate every few months, the importance of studying their impact is undeniable. The broader costs of these events go well beyond the already devastating loss of life and the psychological trauma to survivors.

### v2

@@ 4.1 Mobility Patterns after Mass Shooting
Figure 2(a) plots event study estimates obtained from Model (1) for two distance bands split across median distance from the shooting site. POIs in both distance bands experience a significant decline in weekly visits immediately following the shooting (week 0). Visits then return to pre-event levels and remain at those levels through week 4. This pattern differs from that observed in Figure 1(b) (i). One possible explanation for this discrepancy is that the mobility patterns in Figure 1(b) (i) are computed by aggregating visits across all POIs in the sample, whereas the event study estimates in this section capture per-POI effects. Therefore, the discrepancy may indicate that in the early period, the impact was concentrated on a small number of POIs, possibly those with high visit density, and then spread more broadly across space in later periods, leading to a large decline in both aggregate and per-POI levels.
@@
After week 4, POIs near shooting sites experience sizable, statistically significant declines in weekly visits for the rest of the period. This pattern is robust to an alternative specification using four distance bands instead of two (Online Appendix B.1), where, similar to the patterns in Figure 1(b) (i), the declines in visits within the 1st quartile distance band are small relative to those in the 2nd quartile distance band, which may reflect the presence of authorities, visitors, or the media near the shooting sites.
@@
POIs above the median distance, on the other hand, maintain visits at pre-event levels through week 11 and even exhibit a meaningful increase in visits for much of the remaining period. Two potential factors may explain the increase in weekly visits to these POIs. First, the resumption of previously postponed economic activity by individuals who normally use those POIs. Second, the relocation of activities to further away places by individuals who, prior to the event, typically used the POIs near the shooting site. The former implies only a temporal displacement of activities, but the latter suggests both temporal and spatial displacements.
Given the small initial decline in visits to faraway POIs and then a notable, persistent increase later in the period, the resumption of postponed activities is unlikely to fully explain this rise, suggesting some degree of substitution or relocation of economic activity across regions. That is, visits lost from the POIs in close proximity to the crime scene might have been redistributed, to some extent, to the POIs far from it. We further explore this relocation channel in Section 4.2 by analyzing the impact by POI type and POI substitutability.
@@
This substitution effect, however, does not kick in immediately after the shooting. Instead, several weeks go by before a significant divergence in post-shooting weekly visit trends is observed across regions. This may be due to the difficulty of substitution in the presence of search and switching costs (Klemperer, 1987; Wilson, 2012; Weizsäcker, 1984). Therefore, the relocation of economic activity implies friction costs or efficiency losses due to a deviation from the optimal choices, which, in turn, could affect more macro-level economic outcomes such as wage, unemployment rate, and housing price, among others (a la Brodeur and Yousaf, 2022). [^15] Compared to mere temporal displacement, the relocation of activities can result in more enduring and persistent social costs since the new location is likely sub-optimal compared to the previous one.
@@
Note that the changes in weekly visits in Figure 2(a) are “per POI”; therefore, the aggregate impact can be substantial. Since there are over 150 thousand POIs in the data, the overall impact will be the product of the estimated change per POI and the number of POIs in a given distance band. For instance, the reduction of about 5.4 visits per POI in week 5 for those POIs below the median distance will amount to an overall decrease of over four hundred thousand weekly visits. The median value of nonzero weekly visits to nearby POIs in the dataset during the week prior to the shootings (the baseline week) is 21, and thus a decrease of 5.4 visits constitutes roughly a 26% decline relative to the median.
@@
The differential mobility patterns in the two regions persist over time, resulting in a significant gap in weekly visits. Figure 2(b) plots the DiD estimates for POIs below the median distance (i.e., nearby POIs) with the reference group being the POIs above the median distance (i.e., far away POIs). The graph thus directly quantifies the gap in weekly visits over time between the two regions and its significance. The gap becomes statistically significant after week 4 and increases to about 8 visits in week 17. The same pattern is observed from an alternative specification with four distance bands, although estimates for first-distance-quartile POIs lack precision (see Online Appendix B.1).
@@
Our results in this section are robust to the thresholds used to define distance bands. Online Appendix B.2 provides nonparametric results based on splines (piece-wise polynomials), where we estimate the evolution of weekly visits as a function of “continuous” distance. The patterns are consistent with those reported in this section. The spline results also support setting 5 miles as the cutoff for sample selection, as estimates approach zero.
@@ 4.2.1 Impact by POI Type
Figure 3: Change In Weekly Visits by POI Category
*Note: 95% confidence intervals are shown. Standard errors are clustered at the POI level. “Near”/“Far” refers to POIs below/above median distance from the shooting location. The two numbers shown in parentheses on the right side of the figure correspond to the counts of nearby and far-away POIs, respectively, for each category.
Figure 3 reports the changes in visits and corresponding 95% confidence intervals for various categories in the week immediately after the event (week 0) and approximately four months after the shooting (week 15). [^16] Online Appendix D.1 includes the results for weeks 5 and 10. The estimates indicate that decreases in visits in week 0 are quite general across a variety of POI types in both distance bands, especially closer to the shooting site, implying widespread economic costs of mass shootings.
Notably, for recreation, temporary accommodation, bars and private facilities (e.g., real estate, professional services), the increase in visits to farther-away POIs in week 15 closely mirrors the decrease to nearby POIs, pointing to potential relocation within these categories. Such relocation patterns are observed only for some selected categories, suggesting that POI characteristics may play a role in the substitution effects.
Heterogeneity of the impact across POI types is also notable. Recreation, education, temporary accommodation, food/drink stores exhibit larger changes, although their confidence intervals are also wider. The large and significant effect on education, which follows immediately after shootings and persists over time, deserves a look. For example, the decrease in visits to primary and secondary schools even in week 15 is estimated to be around 15.1 for nearby POIs and 9.3 for faraway POIs, which in aggregate amounts to a reduction of about 24 and 13 thousand visits, respectively. Disruption of educational activities not only hampers student learning, but also has negative effects on parents’ economic activity, such as reduced work hours and productivity (Garcia and Cowan, 2024). Therefore, this large impact on education suggests a substantial and broad social cost.
@@ 4.2.2 Impact by POI Substitutability
The relative size of the changes in weekly visits across different POI types in Figure 3 suggests that ease of substitution might play a role in the observed patterns. It is relatively easy to change the recreational spaces you frequent and thus we see a significant shift in visits from nearby places to far-away places for recreation. In contrast, switching healthcare providers is difficult or undesirable in the short run, which accounts for the negligible effect on healthcare facilities four months after the shooting.
To generalize the idea, we evaluate how the substitutability of POIs is associated with the magnitude of the impact of mass shooting. We perform two DiD regressions, similar to that in Figure 2(b), for two different groups : high-substitutability and low-substitutability POIs. Substitutability is defined in two distinct ways, as shown in Table 4(a). The Common Categories measure of substitutability accounts for the nature of business based on our own experience but can be subjective. The Competition Level measure takes into account the level of competition in a community but may be endogenous since the number of competitors or businesses is determined by local demand or differential preference among consumers. [^17] That is, even though there are many competing locations, it may be a result of many different tastes among consumers, so substitution can still be limited. Each measure has its pros and cons, so the results in this section should be used to gain insights, rather than quantify the difference.
Figures 4(b) and 4(c) report the DiD estimates based on the Common Categories and the Competition Level measures, respectively. In both cases, there are distinct patterns in the evolution of foot traffic over time, with the redistribution of visits from nearby to farther-away POIs being more pronounced for highly substitutable POIs toward the end of the period. Due to the smaller sample size for each DiD regression, these differences are often not statistically significant. Nevertheless, the differential patterns are evident in these plots and the patterns are robust to an alternative specification using 4-quartile distance bands in Online Appendix D.2.
@@ 4.3 Degree of Traumatic Stress and Magnitude of Impact
The previous findings indicate that mass shootings drive significant temporal and spatial displacements of economic activity. Given that the likelihood of another mass shooting occurring at the exact same location is nearly zero, these displacements should be understood in light of the psychological effects mass shootings have on communities. It has been well documented in the psychology and health literatures that traumatic events take a toll not only on survivors but also on their communities. Even those not directly affected by the event can develop post-traumatic stress symptoms, which is especially likely in cases of mass violence (Galea et al., 2003) and for those individuals who are physically closer to the trauma (Furr et al., 2010; Matt and Vasquez, 2008; May and Wisco, 2016).
@@
To examine how the degree of traumatic stress in our context influences the results, we split our sample into high- and low- traumatic stress groups based on three factors: preexisting crime rate, media coverage, and pre-event visit density. We then analyze how these factors affect the magnitude of the impact of mass shooting. Table 5(a) summarizes how we split the data and predictions based on the literature. The analyses here are similar to those in Figures 4(b) and 4(c), and supplemental plots are provided in Online Appendix E.
4.3.1 Preexisting Crime Rates
Two theoretical frameworks from psychology inform our predictions regarding how mass shootings may have different effects depending on the preexisting levels of crime in the affected area. The first is stress inoculation, which states that exposure to smaller instances of stress can promote resilience to future stressful episodes (Ashokan et al., 2016; Gerber et al., 2018). For example, Amir and Sol (1999) find that students in an Israeli university experienced decreased levels of distress when facing a traumatic event if they had previously experienced other traumatic events. The second, desensitization, consists of emotional numbing resulting from repeated exposure to chronic community violence (Kennedy and Ceballo, 2016; Gaylord-Harden et al., 2011; Mrug et al., 2008; Ng-Mak et al., 2004).
It follows from these frameworks that the impact of mass shootings is expected to be greater when it occurs in a previously safer area, where the event would strike people who are less habituated to violent crime. We match each mass shooting with the homicide rate of the county where the shooting occurred in the year prior to the event. Homicide statistics are sourced from County Health Rankings and Roadmaps [^18] https://www.countyhealthrankings.org. Then, we classify the shooting areas into two groups: high-crime areas, defined as those with above-median crime rates, and low-crime areas, including all other locations.
Figure 5(b) plots the evolution of weekly visits for high- and low-crime areas. Over time, the impact of mass shootings on mobility becomes stronger when they occur in previously safer areas. That is, when communities are less habituated to violence, their reaction is stronger, resulting in more significant displacements of activities.
4.3.2 Extent of Media Coverage
Evidence suggests that the psychological impact of crime is driven primarily by perceived risk, not by actual recorded crimes (Pearson and Breetzke, 2014; Braakman, 2012; Hamermesh, 1999; Hale, 1996). The media, then, would play an important role through its influence on public perception. According to Mastrorocco and Minale (2018), who use a staggered introduction of digital TV in Italy as a natural experiment, reduced exposure to crime news channels causes people to decrease their concerns about crime. Similarly, Velasquez et al. (2020) find that spikes in crime reporting increase people’s perceptions about the prevalence of crime, even when actual crime rates decline. It follows that, all else equal, more extensive media coverage on mass shootings is likely associated with greater psychological distress at the community level, and thus stronger impacts on community mobility patterns.
@@ 4.3.2 Extent of Media Coverage
Shooting events are then divided in two groups: those receiving above-median media coverage (high coverage) and those below this threshold (low coverage). Figure 5(c) plots the evolution of weekly visits for high- and low- media coverage shootings. Consistent with our prediction, more extensively publicized events tend to result in larger impacts.
@@ 4.3.3 Visitor Density
Fear of crime is associated with an individual’s perception of vulnerability – that is, the self-assessed likelihood of victimization, which can translate into economically-relevant behavioral changes, such as the avoidance of certain physical areas or situations (Hale, 1996; Garofalo, 1981; Whitley and Prince, 2005). Furthermore, the decision to avoid a physical environment depends on the individual’s assessment of potential risks, even in absence of direct connection between the location and specific instances of crime. (Marquet et al., 2020; Navarrete-Hernandez et al., 2023).
@@
The results are consistent with the prediction that communities respond more intensely for the locations where larger gatherings of people are expected, likely in an effort, conscious or not, to avoid situations that are perceived as conducive to victimization. POIs with low baseline visit density exhibit little to no change in visits after a mass shooting, with very narrow confidence intervals. The median number of nonzero weekly visits prior to the shootings among all POIs in the sample is 20, suggesting that low-baseline-visits POIs are generally very small and, as a result, do not show meaningful variation in visits. [^20] This may indicate that the effects of mass shootings presented elsewhere in the paper can be seen as a lower bound, as they include POIs that are expected to show little change in visits due to the nature of the business.
5 Conclusion